OpenStudy (anonymous):

Find all solutions: 1). x^2+10x-24=0 2). 12m^2-5m=0

OpenStudy (anonymous):

#2) \[12m^2-5m=0\] \[m(12m-5)=0\] \[m=0\] or \[12m-5=0\] \[12m=5\] \[m=\frac{12}{5}\]

OpenStudy (anonymous):

#1) \[x^2+12x-24=0\] \[(x-2)(x+12)=0\] \[x-2=0\] \[x=2\] or \[x+12=0\] \[x=-12\]